# Knowledge Editing for Large Language Models: A Survey

### v3

@@ Abstract.
Large Language Models (LLMs) have recently transformed both the academic and industrial landscapes due to their remarkable capacity to understand, analyze, and generate texts based on their vast knowledge and reasoning ability. Nevertheless, one major drawback of LLMs is their substantial computational cost for pre-training due to their unprecedented amounts of parameters. The disadvantage is exacerbated when new knowledge frequently needs to be introduced into the pre-trained model. Therefore, it is imperative to develop effective and efficient techniques to update pre-trained LLMs. Traditional methods encode new knowledge in pre-trained LLMs through direct fine-tuning. However, naively re-training LLMs can be computationally intensive and risks degenerating valuable pre-trained knowledge irrelevant to the update in the model. Recently, Knowledge-based Model Editing (KME), also known as Knowledge Editing (KE), has attracted increasing attention, which aims to precisely modify the LLMs to incorporate specific knowledge, without negatively influencing other irrelevant knowledge. In this survey, we aim to provide a comprehensive and in-depth overview of recent advances in the field of KME. We first introduce a general formulation of KME to encompass different KME strategies. Afterward, we provide an innovative taxonomy of KME techniques based on how the new knowledge is introduced into pre-trained LLMs, and investigate existing KME strategies while analyzing key insights, advantages, and limitations of methods from each category. Moreover, representative metrics, datasets, and applications of KME are introduced accordingly. Finally, we provide an in-depth analysis regarding the practicality and remaining challenges of KME and suggest promising research directions for further advancement in this field.
@@ 1. Introduction
Recently, large language models (LLMs) have become a heated topic that revolutionizes both academia and industry (OpenAI, 2023; Brown et al., 2020; Touvron et al., 2023; Zhao et al., 2023). With the substantial factual knowledge and reasoning ability gained from pre-training on large corpora, LLMs have exhibited an unprecedented understanding of textual information, which are able to analyze and generate texts akin to human experts (Li and Qiu, 2023; Su et al., 2022; Zhou et al., 2023; Liao and Vaughan, 2023; Song et al., 2023b). Nevertheless, one main drawback of LLMs is the extremely high computational overhead of the training process due to the large amounts of parameters (Ziegler et al., 2019; Honovich et al., 2022; Hu et al., 2023a). This is exacerbated by the continuous evolvement of the world where the requirement of updating pre-trained LLMs to rectify obsolete information or incorporate new knowledge to maintain their relevancy is constantly emerging (Scialom et al., 2022; Luo et al., 2023; Song et al., 2023a; Li et al., 2022). For example, as in Fig. 1, an outdated LLM cannot precisely describe the latest achievements of Lionel Messi, which requires explicit injection of new knowledge to generate the correct answers.
@@
One feasible while straightforward strategy for updating pre-trained LLMs is through naive fine-tuning (Dubois et al., 2023; Taori et al., 2023; Wei et al., 2021; Chung et al., 2022), where parameters of pre-trained LLMs are directly optimized to encode new knowledge from new data (Zhao et al., 2023; Azamfirei et al., 2023; Peng et al., 2023a; Menick et al., 2022). For example, various instruction-tuning-based methods are proposed to fine-tune the pre-trained LLMs on the newly collected corpus in a supervised learning manner (Wang et al., 2022a; Peng et al., 2023b; Wang et al., 2023c; Min et al., 2023). Although such fine-tuning techniques are widely used and capable of injecting new knowledge into LLMs, they are known for the following disadvantages: (1) Even with several parameter-efficient strategies proposed to improve the efficiency (Liu et al., 2022; Zaken et al., 2022; Wang et al., 2022b), fine-tuning LLMs may still require intensive computational resources (Mitchell et al., 2022b; Zheng et al., 2023; Meng et al., 2022). (2) The fine-tuned model could overfit the new data, especially when the dataset utilized for fine-tuning is small in scale (De Cao et al., 2021; Meng et al., 2023; Mitchell et al., 2022a). (3) More importantly, fine-tuning LLMs alters the pre-trained weights with no constraints, which risks losing valuable existing knowledge in LLMs (Madaan et al., 2022; Huang et al., 2023; Dong et al., 2022a). These challenges limit the practicality of fine-tuning techniques in updating LLMs with new knowledge.
Figure 1. An intuitive example of KME for efficient knowledge update of pre-trained LLMs.
@@
To address the drawbacks of naive fine-tuning for updating LLMs, more attention has been devoted to Knowledge-based Model Editing [^1] The concept is also termed as Knowledge Editing or Model Editing. For simplicity, we refer to it as KME in this paper. (KME). In general, KME aims to precisely modify the behavior of pre-trained LLMs to update specific knowledge without negatively influencing other pre-trained knowledge irrelevant to the updates  (Yao et al., 2023; Wang et al., 2023b; Pinter and Elhadad, 2023). In KME, the update of a specific piece of knowledge in LLMs is typically formulated as an edit, such as rectifying the answer to “Who is the president of the USA?” from “Trump” to “Biden”. Regarding a specific edit, KME strategies typically modify the model output by either introducing an auxiliary network (or set of parameters) into the pre-trained model (Li et al., 2023c; Zhong et al., 2023; Hartvigsen et al., 2022) or updating the (partial) parameters to store the new knowledge (Ha et al., 2016; Dai et al., 2022; Li et al., 2023b; Gupta et al., 2023). Through these strategies, the KME techniques can store new knowledge in memories or locate it in model parameters for updating, thereby precisely injecting the knowledge into the model. In addition, certain methods also introduce explicit loss to contain the updating process, such that the edited model maintains consistent behaviors on unmodified knowledge. With these advantages, KME techniques can provide an efficient and effective way to constantly update LLMs with novel knowledge without explicit model re-training.
@@
Although KME shares certain similarities with fine-tuning strategies, it poses unique advantages in updating LLMs, which are worthy of deeper investigations. Particularly, both KME and model fine-tuning seek to update pre-trained LLMs by injecting new knowledge. However, aside from this shared objective, KME focuses more on two crucial properties that cannot be easily addressed by fine-tuning. (1) Locality requires that the edited model does not unintentionally influence the output of other irrelevant inputs with distinct semantics. For example, when the edit regarding the president of the USA is updated, the edited model should not alter its knowledge about the prime minister of the UK. The practicality of KME methods largely relies on their ability to maintain the outputs for unrelated inputs, which serves as a major difference between KME and fine-tuning (Qin et al., 2022). (2) Generality represents whether the edited model can generalize to a broader range of relevant inputs regarding the edited knowledge. Specifically, it indicates the model’s capability to present consistent behavior on inputs that share semantic similarities. For example, when the model is edited regarding the president, the answer to the query of the president’s spouse should also change accordingly. In practice, it is important for KME methods to ensure that the edited model can adapt well to these related input texts. In summary, due to these two unique objectives, KME remains a challenging task that requires specific strategies for satisfactory effectiveness.
@@
Differences between this survey and existing ones. Several surveys have been conducted to examine various aspects of (large) language models (Zhao et al., 2023; Thirunavukarasu et al., 2023; Fan et al., 2023; Kalyan et al., 2022; Chang et al., 2023; Kasneci et al., 2023). Nevertheless, there is still a dearth of thorough investigations that comprehensively encompass the existing literature and the continuous progress in the field of LLM editing. For example, recent works (Wang et al., 2023c; Min et al., 2023) have discussed the fine-tuning strategies to incorporate new knowledge in pre-trained LLMs with more data samples. However, the distinctiveness of KME, i.e., locality and generality, is not adequately discussed, which will be thoroughly analyzed in this survey. Two additional surveys (Fei et al., 2021; Hu et al., 2023b) review knowledge-enhanced language models. However, their main focus is on leveraging external knowledge to enhance the performance of the pre-trained LLMs, without addressing the editing task based on specific knowledge. To the best of our knowledge, the most related paper to our survey is  (Yao et al., 2023), which provides a brief overview of KME and concisely discusses the advantages of KME methods and their challenges. Nevertheless, the investigation lacks a thorough examination of more details of KME, e.g., categorizations, datasets, and applications. Another recent work (Wang et al., 2023b) proposes a framework for KME that unifies several representative methods. This work focuses on the implementation of KME techniques, with less emphasis on the technical details of different strategies. More recently, a work (Pinter and Elhadad, 2023) discusses the limitations of KME methods regarding the faithfulness of edited models, while it is relatively short and lacks a more comprehensive introduction to all existing methods. Considering the rapid advancement of KME techniques, we believe it is imperative to review details of all representative KME methods, summarizing the commonalities while discussing the uniqueness of each method, and discuss open challenges and prospective directions in the domain of KME, which will facilitate the further advancement of the field.
@@
  • In-Depth Analysis. We formulate the task of LLM editing as a constrained optimization problem, where methods from each category can be viewed as a special case with refined constraints. Furthermore, we emphasize the primary insights, advantages, and limitations of each category. Within this context, we delve deep into representative methods from each category and systematically analyze their interconnections.
@@
The remainder of this paper is organized as follows. Section 2 introduces the background knowledge for LLM editing. Section 3 provides a general formulation of the KME task, which can fit into various application scenarios. Section 4 provides a comprehensive summary of evaluation metrics for KME strategies, which is crucial for a fair comparison of various methods. Before delving into the specific methods, we provide a comprehensive categorization of existing methods into three classes in Section 5.1, where their relationship and differences are thoroughly discussed. Then we introduce the methods from the three categories in detail, where the advantages and limitations of each category are summarized. Section 6 introduces the prevalently used public datasets. Section 7 provides a thorough introduction to various realistic tasks that can benefit from KME techniques. Section 8 discusses the potential challenges of KME that have not been addressed by existing techniques. This section also provides several potential directions that can inspire future research. Lastly, we conclude this survey in Section 9.
@@ 2.1. Editing of Machine Learning Models
Model Editing is a special type of model modification strategy where the modification should be as precise as possible. Nevertheless, it should accurately modify the pre-trained model to encode specific knowledge while maximally preserving the existing knowledge, without affecting their behavior on unrelated inputs (Ilharco et al., 2023). First explored in the computer vision field, Bau et al. (2020) investigate the potential of editing generative adversarial networks (GAN) (Goodfellow et al., 2020) by viewing an intermediate layer as a linear memory, which can be manipulated to incorporate novel content. Afterward, Editable Training (Sinitsin et al., 2020) is proposed to encourage fast editing of the trained model in a model-agnostic manner. The goal is to change the model predictions on a subset of inputs corresponding to misclassified objects, without altering the results for other inputs. In (Santurkar et al., 2021), the authors propose a method that allows for the modification of a classifier’s behavior by editing its decision rules, which can be used to correct errors or reduce biases in model predictions. In the field of natural language processing, several works (Dai et al., 2022; Mitchell et al., 2022b) have been proposed to perform editing regarding textual information. Specifically, Zhu et al. (2020) propose a constrained fine-tuning loss to explicitly modify specific factual knowledge in transformer-based models (Vaswani et al., 2017). More recent works (Geva et al., 2021, 2022) discover that the MLP layers in transformers actually act as key–value memories, thereby enabling the editing of specific knowledge within the corresponding layers.
@@
$s$ Subject in a knowledge triple
$r$ Relation in a knowledge triple
$o$ Object in a knowledge triple
@@
$\mathbf{q}^{(l)}_{i}$ Query vector for the $i$-th attention head of the $l$-th attention module in Transformer
$\mathbf{k}^{(l)}_{i}$ Key vector for the $i$-th attention head of the $l$-th attention module in Transformer
$\mathbf{v}^{(l)}_{i}$ Value vector for the $i$-th attention head of the $l$-th attention module in Transformer
@@ 2.2. Large Language Models
Transformers with billions of parameters trained on large corpora have demonstrated emergent ability, showcasing an unprecedented understanding of factual and commonsense knowledge (Zhao et al., 2023). Consequently, these models are referred to as large language models (LLMs) to indicate their drastic distinction from traditional small-scale language models (Thirunavukarasu et al., 2023; Fan et al., 2023). Generally, based on the specific parts of the transformer utilized for language modeling, existing LLMs can be categorized into three classes: encoder-only LLMs, such as BERT (Bidirectional Encoder Representations from Transformers) (Kenton and Toutanova, 2019), encoder-decoder-based LLMs such as T5 (Text-To-Text Transfer Transformer) (Raffel et al., 2020), and decoder-only models (also the most common structures in LLMs) such as different versions of GPT (Generative Pre-trained Transformer) (Radford et al., 2018) and LLaMA (Large Language Model Meta AI) (Touvron et al., 2023).
2.2.3. Training Strategies.
Typically, LLMs with encoder-only structures are trained based on masked language modeling (MLM) (Devlin et al., 2018), which randomly masks out tokens in the corpus and predicts the missing tokens by attending to tokens from both sides. Therefore, token embeddings can be better learned by considering context from a broader scope. On the other hand, LLMs based on encoder-decoder and decoder-only structures utilize the causal language modeling (CLM) strategy for pre-training, which autoregressively predicts the next token by only attending to previous tokens. Therefore, they are more powerful at generating new content based on pretexts.
The training of generative LLMs is mainly based on two stages, i.e., the pre-training stage and the fine-tuning stage. In the first stage, LLMs are trained by language modeling (i.e., next/masked token prediction) on large corpus such as Wikipedia (Vrandečić and Krötzsch, 2014) with commonsense and factual knowledge, ArXiv papers (Beltagy et al., 2019) with scientific/math knowledge, codes from GitHub (Chen et al., 2021) with coding knowledge, etc. Then during the fine-tuning stage, exemplar prompt-answer pairs (e.g., question and answers, demands and codes, description, and formula) are provided to the LLMs such that they are able to answer questions based on the prompt instead of simply completing the sentence. In addition, RLHF (reinforcement learning from human feedback) (Lambert et al., 2022) is usually introduced to further fine-tune the model by learning from human ranks of different generated answers without the explicit requirement of ground truth pairs. After training, the LLM can be viewed as a database, where prompts formed as natural language can be used as queries to obtain knowledge encoded in the weights of transformers (Chang et al., 2023; Kasneci et al., 2023).

### v4

@@ Abstract.
Large Language Models (LLMs) have recently transformed both the academic and industrial landscapes due to their remarkable capacity to understand, analyze, and generate texts based on their vast knowledge and reasoning ability. Nevertheless, one major drawback of LLMs is their substantial computational cost for pre-training due to their unprecedented amounts of parameters. The disadvantage is exacerbated when new knowledge frequently needs to be introduced into the pre-trained model. Therefore, it is imperative to develop effective and efficient techniques to update pre-trained LLMs. Traditional methods encode new knowledge in pre-trained LLMs through direct fine-tuning. However, naively re-training LLMs can be computationally intensive and risks degenerating valuable pre-trained knowledge irrelevant to the update in the model. Recently, Knowledge-based Model Editing (KME), also known as Knowledge Editing or Model Editing, has attracted increasing attention, which aims to precisely modify the LLMs to incorporate specific knowledge, without negatively influencing other irrelevant knowledge. In this survey, we aim to provide a comprehensive and in-depth overview of recent advances in the field of KME. We first introduce a general formulation of KME to encompass different KME strategies. Afterward, we provide an innovative taxonomy of KME techniques based on how the new knowledge is introduced into pre-trained LLMs, and investigate existing KME strategies while analyzing key insights, advantages, and limitations of methods from each category. Moreover, representative metrics, datasets, and applications of KME are introduced accordingly. Finally, we provide an in-depth analysis regarding the practicality and remaining challenges of KME and suggest promising research directions for further advancement in this field.
@@ 1. Introduction
Recently, large language models (LLMs) have become a heated topic that revolutionizes both academia and industry (OpenAI, 2023; Brown et al., 2020; Touvron et al., 2023; Zhao et al., 2023). With the substantial factual knowledge and reasoning ability gained from pre-training on large corpora, LLMs have exhibited an unprecedented understanding of textual information, which are able to analyze and generate texts akin to human experts (Li and Qiu, 2023; Su et al., 2022; Zhou et al., 2023; Liao and Vaughan, 2023; Song et al., 2023b). Nevertheless, one main drawback of LLMs is the extremely high computational overhead of the training process due to the large amounts of parameters (Ziegler et al., 2019; Honovich et al., 2022; Hu et al., 2023a). This is exacerbated by the continuous evolvement of the world where the requirement of updating pre-trained LLMs to rectify obsolete information or incorporate new knowledge to maintain their relevancy is constantly emerging (Scialom et al., 2022; Luo et al., 2023; Song et al., 2023a; Li et al., 2022). For example, as in Fig. 1, the outdated LLM, GPT-3.5, cannot precisely describe the latest achievements of the famous soccer player Lionel Messi, which requires an explicit injection of new knowledge to generate the correct answers.
@@
One feasible while straightforward strategy for updating pre-trained LLMs is through naive fine-tuning (Dubois et al., 2023; Taori et al., 2023; Wei et al., 2021; Chung et al., 2022), where parameters of pre-trained LLMs are directly optimized to encode new knowledge from new data (Zhao et al., 2023; Azamfirei et al., 2023; Peng et al., 2023a; Menick et al., 2022). For example, various instruction-tuning methods are proposed to fine-tune pre-trained LLMs on newly collected data in a supervised learning manner (Wang et al., 2022a; Peng et al., 2023b; Wang et al., 2023d; Min et al., 2023). Although such fine-tuning techniques are widely used and capable of injecting new knowledge into LLMs, they are known for the following disadvantages: (1) Even with parameter-efficient strategies to improve efficiency (Liu et al., 2022; Zaken et al., 2022; Wang et al., 2022b), fine-tuning LLMs may still require intensive computational resources (Mitchell et al., 2022b; Zheng et al., 2023; Meng et al., 2022). (2) Fine-tuning LLMs alters the pre-trained parameters without constraints, which can lead to the overfitting problem, where LLMs face the risk of losing valuable existing knowledge (Zhang et al., 2024).
Figure 1. An example of KME for efficient update of knowledge in LLMs.
@@
To address the drawbacks of updating LLMs with naive fine-tuning, more attention has been devoted to Knowledge-based Model Editing [^1] The concept is also termed as Knowledge Editing or Model Editing. For clarity, we refer to it as KME in this paper. (KME). In general, KME aims to precisely modify the behavior of pre-trained LLMs to update specific knowledge, without negatively influencing other pre-trained knowledge irrelevant to the updates (Yao et al., 2023; Wang et al., 2023c; Pinter and Elhadad, 2023). In KME, the update of a specific piece of knowledge in LLMs is typically formulated as an edit, such as rectifying the answer to “Who is the president of the USA?” from “Trump” to “Biden”. Regarding a specific edit, KME strategies typically modify the model output by either introducing an auxiliary network (or set of parameters) into the pre-trained model (Li et al., 2023b; Zhong et al., 2023; Hartvigsen et al., 2023) or updating the (partial) parameters to store the new knowledge (Ha et al., 2016; Dai et al., 2022; Li et al., 2024c; Gupta et al., 2023). Through these strategies, KME techniques can store new knowledge in new parameters or locate it in model parameters for updating, thereby precisely injecting the knowledge into the model. In addition, certain methods further introduce optimization constraints to ensure that the edited model maintains consistent behaviors on unmodified knowledge (Zhu et al., 2020; Chen et al., 2020; Ni et al., 2023). With these advantages, KME techniques can provide an efficient and effective way to constantly update LLMs with novel knowledge without explicit model re-training (Zhang et al., 2024).
@@
While sharing certain similarities with fine-tuning strategies, KME poses unique advantages in updating LLMs, which are worthy of deeper investigations. Particularly, both KME and model fine-tuning seek to update pre-trained LLMs with new knowledge. However, aside from this shared objective, KME focuses more on two crucial properties that cannot be easily addressed by fine-tuning. (1) Locality requires that KME does not unintentionally influence the output of other irrelevant inputs with distinct semantics. For example, when the edit regarding the president of the USA is updated, KME should not alter its knowledge about the prime minister of the UK. The practicality of KME methods largely relies on their ability to maintain the outputs for unrelated inputs, which serves as a major difference between KME and fine-tuning (Qin et al., 2022). (2) Generality represents whether the edited model can generalize to a broader range of relevant inputs regarding the edited knowledge. Specifically, it indicates the model’s capability to present consistent behavior on inputs that share semantic similarities. For example, when the model is edited regarding the president, the answer to the query about the leader or the head of government should also change accordingly. In practice, it is important for KME methods to ensure that the edited model can adapt well to such related input texts. To summarize, due to these two unique objectives, KME remains a challenging task that requires specific strategies for satisfactory effectiveness.
@@
Differences between this survey and existing ones. Several surveys have been conducted to examine various aspects of (large) language models (Zhao et al., 2023; Kalyan et al., 2022; Thirunavukarasu et al., 2023; Fan et al., 2023; Chang et al., 2023; Kasneci et al., 2023). Nevertheless, there is still a dearth of thorough investigations of existing literature and continuous progress in editing LLMs. For example, recent works (Wang et al., 2023d; Min et al., 2023) have discussed the fine-tuning strategies that inject new knowledge in pre-trained LLMs with more data samples. However, the distinctiveness of KME, i.e., locality and generality, is not adequately discussed, which will be thoroughly analyzed in this survey. Two other surveys (Fei et al., 2021; Hu et al., 2023b) review knowledge-enhanced language models. However, they mainly focus on leveraging external knowledge to enhance the performance of the pre-trained LLMs, without addressing the editing task based on specific knowledge. To the best of our knowledge, the most related work (Yao et al., 2023) to our survey provides a brief overview of KME and concisely discusses the advantages of KME methods and their challenges. Nevertheless, the investigation lacks a thorough examination of more details of KME, e.g., categorizations, datasets, and applications. The following work (Zhang et al., 2024) additionally includes experiments with classic KME methods. Another recent work (Wang et al., 2023c) proposes a framework for KME that unifies several representative methods. This work focuses on the implementation of KME techniques, with less emphasis on the technical details of different strategies. A more recent study (Pinter and Elhadad, 2023) discusses the limitations of KME methods regarding the faithfulness of edited models, while it is relatively short and lacks a more comprehensive introduction to all existing methods. Considering the rapid advancement of KME techniques, we believe it is imperative to review the details of all representative KME methods, summarize the commonalities while discussing the uniqueness of each method, and discuss open challenges and prospective directions in the domain of KME to facilitate further advancement.
@@
  • In-Depth Analysis. We formulate the task of KME as a constrained optimization problem, where methods from each category can be viewed as a special case with refined constraints. Furthermore, we emphasize the primary insights, advantages, and limitations of each category. Within this context, we delve deep into representative methods from each category and systematically analyze their interconnections.
@@
The remainder of this paper is organized as follows. Section 2 introduces the background knowledge for KME. Section 3 provides a general formulation of the KME task, which can fit into various application scenarios. Section 4 provides a comprehensive summary of evaluation metrics for KME strategies, which is crucial for a fair comparison across various methods. Before delving into the specific methods, we provide a comprehensive categorization of existing methods into three classes in Section 5.1, where their relationship and differences are thoroughly discussed. Then we introduce the methods from the three categories in detail, where the advantages and limitations of each category are summarized. Section 6 introduces the prevalently used public datasets. Section 7 provides a thorough introduction to various realistic tasks that can benefit from KME techniques. Section 8 discusses the potential challenges of KME that have not been addressed by existing techniques. This section also provides several potential directions that can inspire future research. Lastly, we conclude this survey in Section 9.
@@ 2.1. Editing of Machine Learning Models
Model Editing is a special type of model modification strategy where the modification should be as precise as possible. Nevertheless, it should accurately modify the pre-trained model to encode specific knowledge while maximally preserving the existing knowledge, without affecting their behavior on unrelated inputs (Ilharco et al., 2023). First explored in the computer vision field, Bau et al. (2020) investigate the potential of editing generative adversarial networks (GAN) (Goodfellow et al., 2020) by viewing an intermediate layer as a linear memory, which can be manipulated to incorporate novel content. Afterward, Editable Training (Sinitsin et al., 2020) is proposed to encourage fast editing of the trained model in a model-agnostic manner. The goal is to change the model predictions on a subset of inputs corresponding to misclassified objects, without altering the results for other inputs. In (Santurkar et al., 2021), the authors propose a method that allows for the modification of a classifier’s behavior by editing its decision rules, which can be used to correct errors or reduce biases in model predictions. In the field of natural language processing, several works (Dai et al., 2022; Mitchell et al., 2022b) have been proposed to perform editing regarding textual information. Specifically, Zhu et al. (2020) propose a constrained fine-tuning loss to explicitly modify specific factual knowledge in transformer-based models (Vaswani et al., 2017). More recent works (Geva et al., 2021, 2022) discover that the MLP layers in transformers actually act as key-value memories, thereby enabling the editing of specific knowledge within the corresponding layers.
@@
$s$/$r$/$o$ Subject/Relation/Object in a knowledge triple
@@
$\mathbf{q}^{(l)}_{i}$/$\mathbf{k}^{(l)}_{i}$/$\mathbf{v}^{(l)}_{i}$ Query/Key/Value vector for the $i$-th head of the $l$-th attention module in Transformer
@@ 2.2. Language Models
Transformers with billions of parameters trained on large corpora have demonstrated emergent ability, showcasing an unprecedented understanding of factual and commonsense knowledge (Zhao et al., 2023). Consequently, these models are referred to as large language models (LLMs) to indicate their drastic distinction from traditional small-scale language models (Thirunavukarasu et al., 2023; Fan et al., 2023). Generally, based on the specific parts of the transformer utilized for language modeling, existing LLMs can be categorized into three classes: encoder-only LLMs, such as BERT (Kenton and Toutanova, 2019), encoder-decoder-based LLMs such as T5 (Raffel et al., 2020), and decoder-only models (also the most common structure in LLMs) such as different versions of GPT (Radford et al., 2018) and LLaMA (Touvron et al., 2023).
2.3. Relevant Topics
KME intersects with several extensively researched topics, yet these techniques cannot effectively address KME-specific challenges (Taori et al., 2023; Wei et al., 2021). The most relevant approach is model fine-tuning (Chung et al., 2022; Azamfirei et al., 2023; Menick et al., 2022), including parameter-efficient fine-tuning (Liu et al., 2022; Zaken et al., 2022; Wang et al., 2022b), which requires fewer parameter updates. However, fine-tuning remains computationally intensive and is often impractical for black-box LLMs (Zhao et al., 2023; Zhang et al., 2024). Another related area is machine unlearning (Nguyen et al., 2022), which aims to remove the influence of individual samples from models. Unlike KME, which focuses on abstract and generalized knowledge updates, machine unlearning targets the elimination of specific training data, making it unsuitable for KME. On the other hand, external memorization KME methods share similarities with RAG (retrieval-augmented generation) (Gao et al., 2023), where a large repository of documents is stored and retrieved as needed to provide contextually relevant information for generating responses. While RAG can introduce new knowledge into LLMs by retrieving recently added documents, it does not effectively update the inherent knowledge within LLMs. Thus, RAG is not suitable for the fundamental knowledge updates that KME seeks to achieve.